The probabilities that Sani , Kato and Titi will hit a target
are 3/4,2/5 and 1/3 respectively.If all the three men shoot once,
what is the probability that the target will be hit only once?

Answer by Edwin McCravy(20065)

The probabilities that Sani , Kato and Titi will hit a target
are 3/4,2/5 and 1/3 respectively.
P(S) = 3/4, P(not S) = 1 - 3/4 = 1/4
P(K) = 2/5, P(not K) = 1 - 2/5 = 3/5
P(T) = 1/3, P(not T) = 1 - 1/3 = 2/3

If all the three men shoot once, what is the probability that the
target will be hit only once?
P(S)×P(not K)×(not T)+P(not S)×P(K)×P(not T)+P(not S)×P(not K)×P(T) =

(3/4)(3/5)(2/3)+(1/4)(2/5)(2/3)+(1/4)(3/5)(1/3) = 5/12

Edwin